What is the strength of the ionic interactions between an ion with 1 positive charges and an ion with 1 negative charges, when the distance between them is 16.4 nanometers and the solvent is benzene, with a dielectric constant of 4.6?


1
Expert's answer
2022-03-08T04:29:02-0500

E = kq1q2 / Dr

k = 9 × 109 N m2/C2

q1, q2 - charges

r - distance between the charges

D - dielectric constant

E = 9 × 109 N m2/C2 × 1 × 1 / (4.6 × 16.4 × 10-9 m) = 1.2 × 1017 N m